OpenStudy (mathmath333):

A person has 3 children with at least one boy. Find the probablitiy of having atleast 2 boys among the children ?

OpenStudy (dan815):

=]

OpenStudy (anonymous):

P(at least 2 boys|1boy)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{A person has 3 children with at least one boy.} \hspace{.33em}\\~\\ & \normalsize \text{Find the probablitiy of having atleast 2 boys } \hspace{.33em}\\~\\ & \normalsize \text{among the children ?} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

how to solve this P(at least 2 boys|1boy) ?

OpenStudy (dan815):

you can click edit and put it in the blue part

OpenStudy (mathmath333):

thnks for suggestion

OpenStudy (anonymous):

P(at least 2 boys | at least 1boy) = P(at least 2 boys and at least 1 boy)/P(at least one boy) P(at least 2 boys)/P(at least one boy)

OpenStudy (mathmath333):

is this like bayes therem

OpenStudy (anonymous):

no just conditional probability Bayes is a little different

OpenStudy (mathmath333):

P(at least 2 boys)/P(at least one boy)=(1/3+1/3)/(3/3)

OpenStudy (anonymous):

not quite... P(At least 2 boys) = P(2 boys) + P(3 boys) P(at least 1 boy) = 1-P(no boys)

OpenStudy (mathmath333):

=(1/3+1/3)/(1-0/3) =?

OpenStudy (anonymous):

P(no boys) = P(all girls) how many children does the couple have? 3. => P(all girls) = P(GGG) = (1/2)(1/2)(1/2) = 1/8

OpenStudy (mathmath333):

final answer=(2/3)/(1/8)=1/12

OpenStudy (dan815):



OpenStudy (dan815):

the area of the smallest circle divded by the bigger circle

OpenStudy (anonymous):

no P(At least 2 boys) = P(2 boys) + P(3 boys) P(3 boys) = P(3 girls) = 1/8 P(2 boys) = 3C2 * (1/8) = 3/8 P(At least 2 boys) = 1/2

OpenStudy (dan815):

Prob of atleast 2 boys = prob of atleast 1 boy - prob of only 1 boy prob of @2boys/prob of @ 1 boy = (prob of atleast 1 boy - prob of only 1 boy)/prob of @ 1 boy = 1- prob_of_1_boy/prob of @ 1 boy = 1- prob of 1 boy/(1-prob of no boys)

OpenStudy (mathmath333):

P(at least 2 boys)/P(at least one boy)=(1/2)/(1/8)=1/4 , is it

OpenStudy (dan815):